OpenStudy (gabbyalicorn):

What is the MAD of the data set? {5, 12, 1, 6, 11} mean is 7

OpenStudy (gabbyalicorn):

this part im stuck on :/ is it like /5-7/ and etc @freeG13

OpenStudy (freeg13):

add all of them up then divide them by 5 because that is how numbers there are

OpenStudy (gabbyalicorn):

i did that i got 7 for the mean l5-7l =2 l12-1l=11 l1-7l=6 l6-7l=1 l11-7l=4

OpenStudy (freeg13):

"how many numbers there are"

OpenStudy (gabbyalicorn):

5

OpenStudy (freeg13):

brb

OpenStudy (gabbyalicorn):

i did 5, +12, +1,+ 6,+ 11= 35 divided by 5

OpenStudy (gabbyalicorn):

im trying to do MAD

OpenStudy (freeg13):

what is Mad again?

OpenStudy (gabbyalicorn):

the mean absolute deviation

OpenStudy (rainbow_rocks03):

what do you think the answer is @gabbyalicorn

OpenStudy (freeg13):

The steps to find the MAD include: 1. find the mean (average) 2. find the difference between each data value and the mean 3. take the absolute value of each difference 4. find the mean (average) of these differences

OpenStudy (gabbyalicorn):

i dont know 0.96 thats what i got when i did the mad

OpenStudy (gabbyalicorn):

i tried

OpenStudy (gabbyalicorn):

let me retry

OpenStudy (gabbyalicorn):

k so 1. is 7

OpenStudy (gabbyalicorn):

2. im still stuck on

OpenStudy (freeg13):

1.6

OpenStudy (gabbyalicorn):

howd u get that?

OpenStudy (freeg13):

I looked it up and it is either 1.6 or 3.6

OpenStudy (gabbyalicorn):

you looked up the problem -_-

OpenStudy (gabbyalicorn):

but what if its wrond 0.0

OpenStudy (gabbyalicorn):

wrong*

OpenStudy (freeg13):

sorry I never did MAD i only did mean

OpenStudy (gabbyalicorn):

oh but i got 7 for the mean

OpenStudy (freeg13):

ok well bye

OpenStudy (gabbyalicorn):

:/

OpenStudy (freeg13):

confusing :( :/

OpenStudy (gabbyalicorn):

@Mehek14 @geerky42

OpenStudy (gabbyalicorn):

@YoungTrilogyFT3

OpenStudy (gabbyalicorn):

@satellite73

Mehek (mehek14):

the mean: 7 now tell me 7-5 = 12-7= 7-1= 7-6= 11-7=

OpenStudy (gabbyalicorn):

1. 2 2. 5 3. 6 4. 1 5. 4

OpenStudy (gabbyalicorn):

so 2 + 5 + 6 + 1 + 4 = 3.6 is this correct? @Mehek14

OpenStudy (gabbyalicorn):

someone please help me :''(

Mehek (mehek14):

yes 3.6 is the MAD